The front 1.20 m of a 1,500-kg car is designed as a "crumple zone" that collapses to absorb the shock of a collision.
(a) If a car traveling 27.0 m/s stops uniformly in 1.20 m, how long does the collision last?
s

(b) What is the magnitude of the average force on the car?
N

(c) What is the magnitude of the acceleration of the car? Express the acceleration as a multiple of the acceleration of gravity.
g
Answer #1

Ou 0.11y ings, rea au:
Answer #2
Q:
The front of a 1500 kg car is designed to absorb the shock of a collision by having a "crumple zone" in which the front 1.20 m of the car collapses in absorbing the shock of a collision. If a car traveling 21.0 m/s stops uniformly in 1.20 m,

(a) how long does the collision last,
__________s
(b) what is the magnitude of the average force on the car, and
__________N
(c) what is the acceleration of the car in g's?
__________g's




ANS:
a) use formula s=0.5(v+u)t. 1.2=0.5x21x t. Rearrange to get t=0.114s

b)force=change in momentum/time. force is 1500x21/0.114=275625N

c)To find a use v=u+at. 0=21+0.114a rearrange to find a=-183.75. In g's 183.75/9.8=18.75g
Answer #3
a) use formula s=0.5(v+u)t. 1.2=0.5x27x t. Rearrange to get t=0.08s
b)force=change in momentum/time. force is 1500x27/0.08=506250
c)To find a use v=u+at. 0=27+0.08a rearrange to find a=-337.5. In g's 337.5/9.8=34.43g